# Towards Efficient Transferable Preemptive Adversarial Defense

### v3

@@ VII-B Backbone Model
Figure 7: Different backbone models learn different features. PreActResNet-18 [5], RaWideResNet-70-16 [7], and Bi-Level [10] were adversarially trained, so they learned more crucial features, i.e., ones that are visually similar to the objective.
TABLE VII: Performance of different backbone models.
Ablation Time Cost (s) White-box Defense Transferable Defense
11
Backbone Clean (%) PGD (%) AutoAttack (%) Clean (%) PGD (%) AutoAttack (%) Clean (%) PGD (%) AutoAttack (%)
Bi-Level [10] 89.1 43.9 42.4 28 95.5 84.9 81.8 95.3 72.7 70.8
RaWideResNet-70-16 [7] 93.5 74.5 71.4 178 95.7 91.6 89.9 95.2 78.1 76.2
PreActResNet-18 [5] 88.1 61.8 58.6 26 93.8 85.9 83.3 95.4 79.6 77.3
@@
  • To evaluate the performance of the backbone models in our framework, we followed the results from the previous section and used ResNeXt29-32x4d [37] as the classifier model, and 100 steps backward propagation as Bi-Level [10]. The transferable defenses were all evaluated using XCiT-L12 [6] as the victim model, different from the backbone models.
@@
As shown in Table VII, the best adversarially trained model was RaWideResNet-70-16 [7], so it is not surprising that its white-box defense outperformed that of the other backbone models. However, this most robust model failed to guarantee the best transferability, and the time cost was much higher due to its complexity. We thus empirically selected PreActResNet-18 [5] as the backbone model in Fast Preemption as it had the least time cost and best transferability. Nevertheless, it is worth investigating more comprehensive approaches for determining the backbone model in future work.
VII-C Forward-Backward (Fore-Back) Cascade Learning
@@
As observed in Figure 8, forward (F) propagation accelerated defense but led to overfitting (e.g., F100-B0). In contrast, backward (B) propagation required more steps to converge (e.g., F0-B100) but alleviated overfitting by the forward propagation (e.g., F10-B90). Additionally, as shown by the curve for “Alternate,” the number of steps in forward propagation must be smaller than that in backward propagation to alleviate overfitting.
@@
Our goal is to reduce the time cost, i.e., learning steps, so we further evaluated forward and backward propagation by limiting the learning steps, as illustrated in Figure 9. The results show that when the number of learning steps is small, the combination of forward and backward propagation outperforms backward propagation alone (e.g., F1-B2 vs. F0-B3). Therefore, Fast Preemption runs one step forward and two steps backward propagation learning as the optimal strategy.
@@
VII-D Optimization Algorithm
The results in Table VIII indicate that averaging $N$ FGSM examples results in the best performance with the least time cost. Using PGD results in much more time being spent for preemptive defense, whereas using FGSM but without averaging leads to poor performance.
TABLE VIII: Performance for different $N$ and optimization algorithms.
@@ VII-D Optimization Algorithm
  • Forward propagation ran one step and backward propagation ran two steps. Transferable defense was evaluated using RaWideResNet-70-16 [7] as the victim model.
VIII Conclusion
We have presented an efficient transferable preemptive adversarial defense that we call Fast Preemption. We demonstrated that with its framework and Fore-Back cascade learning algorithm, Fast Preemption offers transferable protection against adversarial attacks among unknown victim models without requiring ground truths. In terms of the time cost, clean accuracy, and robust accuracy, Fast Preemption outperformed benchmark training-time, test-time, and preemptive defenses. We also demonstrated that the protection from Fast Preemption is irreversible unless the backbone model, algorithm, and settings of the defense are all compromised.
This paper lays the theoretical foundation for a preemptive defense framework that can shape future research on each component’s optimization instead of end-to-end application in the relevant field. The toolbox that will be released supports further development and offers a platform for discussing relevant technologies.

### v4

@@ Appendix F Ablation Studies
Figure 7: Different backbone models learn different features. PreActResNet-18 [17], WideResNet-94-16 [5], and Bi-Level [29] were adversarially trained, so they learned more crucial features, i.e., ones that are visually similar to the objective.
F.1 Classifier
In all proactive defenses, correct labeling guarantees that optimization moves in the correct direction. As demonstrated in Tab. 5, a classifier with higher accuracy leads to better performance. As this classifier is not involved in the Fore-Back learning, the time cost of labeling inputs can be ignored compared with that of Fore-Back learning. Note that the use of ground truths is ideal, but this requires manual labeling, which is impractical for efficient user interaction. Therefore, we used the best pretrained classifier in the Fast Preemption framework, one that differs from the backbone model.
Table 6: Performance of different backbone models.
Backbone Model: Bi-Level [29] WideResNet-94-16 [5] PreActResNet-18 [17]
Backbone Robustness Clean (%) 89.1 93.9 88.1
PGD (%) 43.9 77.1 61.8
AutoAttack (%) 42.4 74.0 58.6
Time Cost (s) for Defense 28 69 26
White-Box Defense Clean (%) 95.5 95.8 93.8
PGD (%) 84.9 91.6 85.9
AutoAttack (%) 81.8 90.2 83.3
Transferable Defense Victim: RaWideResNet-70-16 [32] Clean (%) 95.4 95.5 95.3
PGD (%) 85.7 89.7 87.6
AutoAttack (%) 82.7 88.0 85.9
Transferable Defense Victim: XCiT-L12 [11] Clean (%) 95.3 95.3 95.4
PGD (%) 72.7 78.6 79.6
AutoAttack (%) 70.8 75.7 77.3
@@ F.1 Classifier
  • To evaluate the performance of the backbone models in our framework, we followed the results from the previous section and used ResNeXt29-32x4d [19] as the classifier model, and 100 steps backward propagation as Bi-Level [29]. The transferable defenses were all evaluated using RaWideResNet-70-16 [32] and XCiT-L12 [11] as the victim models, different from the backbone models. Bold indicates the best performance.
F.2 Backbone Model
The backbone model determines the defense time cost and performance. As illustrated in Fig. 7, different backbone models learn different features. To evaluate the effect of the backbone model, we also evaluated white-box defenses, for which the backbone model is the same as the victim model, i.e., paired modules, as stated in Sec. 1, although we believe that this setting is not secure due to potential privacy leakage due to the existence of a backdoor [35].
@@ F.2 Backbone Model
As shown in Tab. 6, the best adversarially trained model was WideResNet-94-16 [5], so it is not surprising that its white-box defense outperformed that of the other backbone models. However, this most robust model failed to guarantee the best transferability (see results against XCiT-L12 [11]), and the time cost was much higher due to its complexity. We thus empirically selected PreActResNet-18 [17] as the backbone model in Fast Preemption as it had the least time cost and comparable transferability compared with the SOTA robust model [5]. Nevertheless, it is worth investigating more comprehensive approaches for determining the backbone model in future work.
F.3 Forward-Backward Cascade Learning
@@
As observed in Fig. 8, forward (F) propagation accelerated defense but led to overfitting (e.g., F100-B0). In contrast, backward (B) propagation required more steps to converge (e.g., F0-B100) but alleviated overfitting by the forward propagation (e.g., F10-B90). Additionally, as shown by the curve for “Alternate,” the number of steps in forward propagation must be smaller than that in backward propagation to alleviate overfitting.
@@
Our goal is to reduce the time cost, i.e., learning steps, so we further evaluated forward and backward propagation by limiting the learning steps, as illustrated in Fig. 9. The results show that when the number of learning steps is small, the combination of forward and backward propagation outperforms backward propagation alone (e.g., F1-B2 vs. F0-B3). Therefore, Fast Preemption runs one step forward and two steps backward propagation learning as the optimal strategy.
@@
Table 7: Performance for different $N$ and optimization algorithms.
@@
  • Forward propagation ran one step and backward propagation ran two steps. Transferable defense was evaluated using RaWideResNet-70-16 [32] as the victim model. Bold indicates the best performance.
F.4 Optimization Algorithm
The results in Tab. 7 indicate that averaging $N$ FGSM examples results in the best performance with the least time cost. Using PGD results in much more time being spent for preemptive defense, whereas using FGSM but without averaging leads to poor performance.